OpenStudy (anonymous):

The sum of twice a number and 4 less than than the number is the same difference between -32 and the number. What is the number? I have an equation 2n + (n-4) = n - (-32) but not sure how to solve.

OpenStudy (zzr0ck3r):

2n+(n-4) = n-(-32) 2n+n-4 = n+32

OpenStudy (zzr0ck3r):

can you go from there?

OpenStudy (zzr0ck3r):

first subtract n from both sides and add 4 to both sides and tell me what you get

OpenStudy (anonymous):

2 = 36

OpenStudy (zzr0ck3r):

2n+n-4-n=n+32-n 2n-4 = 32 2n-4+4 = 32+4 2n= 36 you lost the n somehow

OpenStudy (anonymous):

then i take 36/2?

OpenStudy (zzr0ck3r):

correct

OpenStudy (anonymous):

so my answer would be 18